Grade 11Mechanics

in a x ray machine an electron is subjected to a force of 10-23 . in how much time the electron will cover a distance of 0.1 m ? take mass of elec- 10-3 kg

suraj
6 Years agoGrade 11

1 Answer

Arun
6 Years ago

F = ma = 10-23 N

here,

m = 9.1 x 10-31 kg

so, acceleration produced

a = F/m = 10-23 N /  9.1 x 10-31

thus,

a = 1.098 x 107 m/s2

now,

s = ut + (1/2)at2

if electron starts from rest, u = 0

s = (1/2)at2

or

time taken

t = (2s/a)1/2 = (2x0.1 / 1.098 x 107 )1/2

thus, we get

t = 1.35 x 10-4 s